2. Consider the following scores on a mid-term exam.

Emma

60

Noah

75

Olivia

95

Mason

80

Liam

65

Sophia

82

Isabella

75

Ethan

98

James

93

Charlotte

87

2a. What is the mean of the mid-term test scores?

2b. What is the standard deviation of the mid-term test scores?

2c. What is the coefficient of variance for the mid-term tests scores?  

2. Consider the following scores on a mid-term exam.

Emma

60

Noah

75

Olivia

95

Mason

80

Liam

65

Sophia

82

Isabella

75

Ethan

98

James

93

Charlotte

87

2a. What is the mean of the mid-term test scores?

2b. What is the standard deviation of the mid-term test scores?

2c. What is the coefficient of variance for the mid-term tests scores?  
Answer #1

2)

S.No X1 X12
1 60.000 3600.0000
2 75.000 5625.0000
3 95.000 9025.0000
4 80.000 6400.0000
5 65.000 4225.0000
6 82.000 6724.0000
7 75.000 5625.0000
8 98.000 9604.0000
9 93.000 8649.0000
10 87.000 7569.0000
ΣX 810 ΣX2 67046.0000
=Σx/n= 81.000 s12=(Σx2-(Σx)2/n)/(n-1)= 159.556
s1=√s12                                = 12.632
Coefficient of variation = s/x = 15.59%

from above :a )mean =81.00

b) std deviation =12.632

c) coefficient of varation =(s/x)*100 =15. 59%